# Black Hole from Entropy Maximization

### v4

@@ Abstract
One quantum characterization of a black hole motivated by (local) holography and thermodynamics is that it maximizes thermodynamic entropy for a given surface area. In the context of quantum gravity, this could be more fundamental than the classical characterization by a horizon. As a step, we explore this possibility by solving the 4D semi-classical Einstein equation with many matter fields, and find a picture of a black hole. For spherical static highly-excited configurations, we apply local typicality and estimate the entropy including self-gravity to derive its upper bound. The saturation condition uniquely determines the entropy-maximized configuration: self-gravitating quanta condensate into a radially-uniform dense configuration with no horizon, where the self-gravity and a large quantum pressure induced by the curvatures are balanced and no singularity appears. The interior metric is a self-consistent and non-perturbative solution for Planck’s constant. The maximum entropy, given by the volume integral of the entropy density, agrees with the Bekenstein-Hawking formula through self-gravity, deriving the Bousso bound for thermodynamic entropy. Finally, 10 future prospects are discussed, leading to the speculative view that the configuration represents semi-classically a quantum-gravitational condensate with holographic bulk dynamics.
@@ I Introduction
One quantum characterization of a black hole is that it maximizes thermodynamic entropy Dvali3 ; Oriti1 . This is not yet fully understood, but is motivated by a variety of facts. First, the microscopic origin of thermodynamic entropy is quantum: $S=\log\Omega$, where $\Omega$ is the number of quantum states $\{|\psi\rangle\}$ consistent with fixed macroscopic quantities. Therefore, the characterization by entropy is valid in a fully quantum context. Second, gravity is universal in that anything with energy attracts each other. In a strong gravity limit, any spherical configuration with mass $\frac{R}{2G}$ will collapse to a black hole with size $R$. This implies that a black hole can be considered as a macroscopic state with maximum entropy according to the second law of thermodynamics. Third, in the case of uncharged spherical symmetry, the Bekenstein-Hawking formula saturates the Bousso bound $S\leq\frac{{\cal A}}{4l_{p}^{2}}$ Bousso1 : a conjecture that the entropy $S$ inside a finite region is bounded by the boundary surface area ${\cal A}$, which has proposed holographic principle Bousso2 ($l_{p}\equiv\sqrt{\hbar G}$).111 Note here that the value of the Bekenstein-Hawking formula is determined by conserved charges such as ADM energy and does not depend on the presence of a horizon. See also Sec.VII. From these, the entropy maximization should be a candidate quantum characterization of black holes, eventually leading to a quantum description without classical geometry.
Then, what macroscopic quantities should we fix for entropy of a self-gravitating system? When calculating entropy of a non-gravitating object microcanonically, one fixes the energy and volume and counts up quantum states consistent with them. In a self-gravitating system, however, energy and volume are related to each other by the Einstein equation, and the two cannot be fixed independently. Rather, motivated by (local) holography Bousso2 ; Laurent , it should be natural to fix the surface area ${\cal A}$ of the boundary of a finite region, consider the entropy of the interior, and find the maximizing configuration Dvali3 ; Oriti1 .
@@
Thus, as a first step toward a full quantum description, we consider spherically-symmetric static configurations and explore the characterization that a black hole maximizes thermodynamic entropy for a given surface area ${\cal A}=4\pi R^{2}$ $(R\gg l_{p})$, by using the 4D semi-classical Einstein equation, the self-consistent equation in a mean-field approximation of quantum gravity BD ; Kiefer :
@@
where gravity is described by classical metric $g_{\mu\nu}$ and matter by quantum operators. We then construct the interior metric $g_{\mu\nu}^{*}$ that maximizes the entropy, and find a candidate picture of quantum black holes. (Here, $X^{*}$ or $X_{*}$ denotes a quantity $X$ for the entropy-maximized configuration.) Note that we do not assume a priori that a black hole has a horizon or that the maximum entropy is given by the Bekenstein-Hawking formula.
@@
Solving the saturation condition for the entropy bound under the maximum excitation and using the consistency with local thermodynamics, we find the entropy-maximized configuration uniquely (Sec.IV). The interior metric $g_{\mu\nu}^{*}$ is given by
@@
We call it a semi-classical gravity condensate. Here, the near-Planckian curvature inside induces quantum fluctuations of various modes, which generate a large tangential pressure $\langle T^{\theta}{}_{\theta}\rangle$ associated with the 4D Weyl anomaly, making the condensate locally anisotropic. This pressure supports the system against the strong self-gravity, exceeds the Buchdahl limit Buchdahl , and self-consistently keeps the curvature finite. As a result, the energy is distributed throughout the interior, and the small central region ($0\leq r\lesssim\sqrt{\sigma}$) beyond the description by (1) has only a small energy and can be assumed to be almost flat. Thus, no singularity appears.
@@
This is close to but still outside $r=a_{0}$, and therefore, the configuration has no horizon but looks like a classical black hole from the outside. Indeed, it is almost black due to the exponentially large redshift of (3).
We can also obtain the gravity condensate by considering a formation process. Generically, for a given system, a configuration obtained by a reversible process should be thermodynamically typical and have the maximum entropy. In Refs.KMY ; KY1 ; KY4 ; KY5 , we considered a process in which thermal radiation comes together reversibly due to self-gravity in a heat bath at Hawking temperature, solved the self-consistent time evolution including the backreaction from Hawking-like radiation during the formation process, and obtained the metric (3). Therefore, the gravity condensate should be the most typical configuration with the maximum entropy according to the second law of thermodynamics, which is consistent with the above construction based on typicality.
Now, to evaluate explicitly $S_{max}\equiv S[g_{\mu\nu}^{*};R)$, we apply the Unruh effect (or the local temperature due to the particle creation inside) and thermodynamic relations locally to the interior metic (3) and obtain the entropy density $s_{*}(r)$ (Sec.V). We evaluate (2) and derive
@@
where ${\cal A}\equiv 4\pi R^{2}=4\pi a_{0}^{2}+\mathcal{O}(1)$ for (4). Therefore, the maximum entropy $S_{max}$ coincides with the Bekenstein-Hawking formula. Here, the self-gravity changes the entropy (2) from the volume law to the area law Y1 .
This derives the Bousso bound for thermodynamic entropy (Sec.VI). Applying $\nabla_{\mu}\tilde{s}^{\mu}=0$ and a regularity condition to all the configurations in our class, we can show that (2) agrees with one evaluated on a light-sheet $\Sigma$ introduced in Ref.Bousso1 and thus is covariant. Therefore, (5) means that the gravity condensate is derived as the unique configuration saturating the Bousso bound. Furthermore, the interior metric (3) saturates the local sufficient conditions for the Bousso bound proposed in Refs.FMW ; BFM and thus has a holographic bulk dynamics.
We finally discuss 10 prospects for this picture of black holes (Sec.VII): role of self-gravity in holography, relation to other gravity-condensate models Dvali3 ; Oriti1 , gravity-condensate phase/state, path-integral evaluation of $S[g_{\mu\nu};R)$ BY1 ; BY2 ; Jacobson_vol , thermodynamic entropy vs entanglement entropy Minic ; Casini ; Jacobson_entangle , recovery of state-dependence in Hawking radiation KY2 , a picture of non-typical black holes, relation to the classical picture of black holes, gravitational field with finite entropy Jacobson ; Pad , and phenomenology echo ; CY . We then reach the speculative view that the gravity condensate represents semi-classically a mixture of gravity quanta and matter quanta.
@@ II.2 Local typicality
In general, the order of the magnitude of the entropy in a system with a small volume $\Delta V$ and a high temperature $T$ can be estimated as $\sim\langle-T^{t}{}_{t}\rangle_{th}\Delta V/T$. For example, in a ultra-relativistic spherical fluid with volume $4\pi r^{2}\Delta\hat{r}$ (including self-gravity), the Stefan-Boltzmann law holds approximately Landau_SM ; fluidbook , and the entropy is given by $s(r)\Delta\hat{r}=\frac{4\langle-T^{t}{}_{t}(r)\rangle_{th}}{3T_{loc}(r)}4\pi r^{2}\Delta\hat{r}$ (see (76)). Note here that the non-locality of entropy BFM ; BCFM is considered for the width (8). Therefore, using this, (9) and (10), the order of the magnitude of the entropy in our subsystem can be estimated by $s(r)\Delta\hat{r}\sim\frac{\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle}{\epsilon(r)}4\pi r^{2}\Delta\hat{r}$, leading to [^8] One might think that in a pure state, entropy is zero. Microscopically, thermodynamic entropy is given by the number of possible states consistent with fixed macroscopic parameters (most such states are typical), which is state-independent Landau_SM . Physically, for example, a cold atomic system in a pure state will develop thermal behavior after a quench process, such that macroscopic quantities such as energy density have the same value as their thermal expectation values. Thus, assuming local typicality, the typical (i.e., thermal) behavior of the energy density can be used to estimate the entropy density in this phenomenological method. (See Sec.VII for microscopic methods.)
@@ II.3 Self-gravity
which provides the entropy $S[g_{\mu\nu};R)$ for the self-consistent solution to (1). Thus, the Hamiltonian constraint (13) relates one metric $g_{\mu\nu}$ and a set of the typical states $\{|\psi\rangle\}$ that have the same thermodynamic energy density $\langle-T^{t}{}_{t}(r)\rangle_{th}$, and the finite entropy (16) is obtained. [^9] Here are some comments on the subtleties of the above entropy evaluation.We use only the geometrical static condition and local consistencies with thermodynamics, and do not assume a priori a global thermodynamic equilibrium. This is motivated by two facts. More physically, even for static configurations, one should consider their formation processes. Generically, mechanical and global/local thermodynamic equilibria are different due to differences in relaxation time scales Landau_SM . In particular, we are now considering self-gravity. Therefore, it is necessary to consider time-delay effects in the formation process and discuss in which equilibrium state the configuration is for the time scale under consideration. The other is that, a notion of global thermodynamic equilibrium in self-gravitating systems, leading often to thermodynamical instability from negative specific heat Antonov ; Lynden , is still controversial Landau_SM ; Pad_thermo , and in particular, one consistent with (1) is not known. Furthermore, we do not assume a condition of isotropic fluid Green ; Xia (like thermal radiation Sorkin ) because it is not clear a priori whether such fluid maximizes entropy including self-gravity dynamically. Thus, we have estimated $S[g_{\mu\nu};R)$, (16), by using only the geometric static condition and the local consistency with thermodynamics. For the case of $S_{max}$, we will check the validity of this treatment by constructing the self-consistent solution $(g_{\mu\nu}^{*},|\psi\rangle_{*})$ of (1) (Sec.IV.2). In Sec.VII we will discuss its global equilibrium briefly, and in Appendix D we will argue that Tolman’s law does not lead to maximum entropy.
@@ III Upper bound
in the interior metric of (7): a quantum that constitutes a part at $r$ must be at least its wavelength out of the Schwarzschild radius $a(r)$ of the energy inside it Sorkin . (A dynamical origin of (18) will be discussed in Sec.IV.3.1.) Using (11) and $\lambda\sim\frac{\hbar}{\epsilon}$, we can rewrite (18) as
@@
where we have used $4\pi\int^{R}_{0}drr^{3}\langle-T^{t}{}_{t}(r)\rangle\leq 4\pi R\int^{R}_{0}drr^{2}\langle-T^{t}{}_{t}(r)\rangle=R\frac{a(R)}{2G}\approx R\frac{a_{0}}{2G}\equiv RM_{0}$ from (13). [^10] In Ref.Sorkin , (III) and (22) were obtained only for thermal radiation, while our derivation can be applied to more general cases. Note that $G$ doesn’t appear here, but this is a result of the dynamics of gravity because we need $\mathcal{H}=0$ (13) to obtain the ADM energy $M_{0}$.
@@ IV.1 Saturating energy distribution $a_{*}(r)$
The estimation (16) and bound (17) mean that the maximum entropy is obtained by setting $\epsilon_{*}(r)=\epsilon_{max}$. Therefore, (23) becomes $nl_{p}^{2}\sim r(r-a_{*}(r))$, leading to
@@
with $\sigma=fnl_{p}^{2}$ ($f>0$: a constant of $\mathcal{O}(1)$). Thus, the maximum entropy $S_{max}\equiv S[g_{\mu\nu}^{*},R)$ is estimated from (21) and (24) as
@@
where $a_{0}\approx a_{*}(R)\approx R$. (In Sec.V.2, the relation of $a_{0}$ and $R$ will be determined explicitly.) Note that the saturation of both (17) and (18) corresponds to the strong-gravity limit.
@@
where we have applied $\epsilon(r)=\epsilon_{max}$ and (24) to (14). This means that $\mathcal{O}(\sqrt{n})$ bit of information is packed per the Planck length constantly KY2 ; Y1 . Thus, a necessary condition for the entropy maximization is the radial uniformity. Conversely, as a result of radial uniformity, (24) can be uniquely obtained as a solution satisfying the entropy maximization condition (23), without using $\epsilon(r)=\epsilon_{max}$ (see Appendix C). Therefore, from these two arguments, a necessary and sufficient condition for the entropy maximization is radial uniformity. This is a non-trivial result for a self-gravitating system.
@@
On the other hand, using (13) and (24), we have the energy density $\langle-T^{t}{}_{t}\rangle_{*}\approx\frac{1}{8\pi Gr^{2}}$ and the local energy $\Delta E_{loc}^{*}(r)\equiv 4\pi r^{2}\langle-T^{t}{}_{t}\rangle_{*}\Delta\hat{r}_{*}\sim\sqrt{n}m_{p}$. Therefore, we must have at least $\frac{a_{*}(r)}{2G}\gtrsim\sqrt{n}m_{p}$, meaning that (24) holds only in

### v5

@@ Abstract
One quantum characterization of a black hole motivated by (local) holography and thermodynamics is that it maximizes thermodynamic entropy for a given surface area. In the context of quantum gravity, this could be more fundamental than the classical characterization by a horizon. As a step, we explore this possibility by solving the 4D semi-classical Einstein equation with many matter fields. For highly-excited spherically-symmetric static configurations, we apply local typicality and estimate the entropy including self-gravity to derive its upper bound. The saturation condition uniquely determines the entropy-maximized configuration: self-gravitating quanta condensate into a radially-uniform dense configuration with no horizon, where the self-gravity and a large quantum pressure induced by the curvatures are balanced and no singularity appears. The interior metric is a self-consistent and non-perturbative solution in Planck’s constant. The maximum entropy, given by the volume integral of the entropy density, agrees with the Bekenstein-Hawking formula through self-gravity, deriving the Bousso bound for thermodynamic entropy. Finally, 10 future prospects are discussed, leading to a speculative view that the configuration represents a quantum-gravitational condensate in a semi-classical manner.
@@ I Introduction
One quantum characterization of a black hole is that it maximizes thermodynamic entropy Dvali3 ; Oriti1 . This is not yet fully understood, but it can be motivated by various facts. First, the microscopic origin of thermodynamic entropy is quantum: $S=\log\Omega$, where $\Omega$ is the number of quantum states $\{|\psi\rangle\}$ consistent with fixed macroscopic quantities. Therefore, the characterization by entropy is valid in a fully quantum context. Second, gravity is universal in that anything with energy attracts each other. In a strong gravity limit, any spherical configuration with mass $\frac{R}{2G}$ will collapse to a black hole with size $R$. This implies that a black hole can be considered as a macroscopic state with maximum entropy according to the second law of thermodynamics. Third, in the case of uncharged spherical symmetry, the Bekenstein-Hawking formula saturates the Bousso bound $S\leq\frac{{\cal A}}{4l_{p}^{2}}$ Bousso1 : a conjecture that the entropy $S$ inside a finite region is bounded by the boundary surface area ${\cal A}$, which has proposed holographic principle Bousso2 ($l_{p}\equiv\sqrt{\hbar G}$).111 Note here that the value of the Bekenstein-Hawking formula is determined by conserved charges such as ADM energy and does not depend on the presence of a horizon. See also Sec.VII.
Then, what macroscopic quantities should we fix for entropy of a self-gravitating system? When calculating entropy of a non-gravitating object microcanonically, one fixes the energy and volume and counts up quantum states consistent with them. In a self-gravitating system, however, energy and volume are related to each other by the Einstein equation, and the two cannot be fixed independently. Rather, motivated by (local) holography Bousso2 ; Laurent , it should be natural to fix the surface area ${\cal A}$ of the boundary of a finite region and consider the configuration that maximizes the entropy Dvali3 ; Oriti1 .
@@
These suggest a candidate for the quantum definition of black holes: A black hole is the configuration that maximizes the thermodynamic entropy for a fixed surface area ${\cal A}$. This is simple and generic in the sense that it depends only on thermodynamic entropy and surface area, and such a configuration can be found in certain quantum-gravity models that have these two notions at some level.
In this paper, as a first step, we explore this possibility for spherically-symmetric static configurations with ${\cal A}=4\pi R^{2}$ in the 4D semi-classical Einstein equation:
@@
the self-consistent equation in a mean-field approximation of quantum gravity where gravity is described by classical metric $g_{\mu\nu}$ and matter by quantum operators BD ; Kiefer (here, $R\gg l_{p}$). We then construct the interior metric $g_{\mu\nu}^{*}$ that maximizes the entropy, and find a candidate picture of quantum black holes. (Here, $X^{*}$ or $X_{*}$ denotes a quantity $X$ for the entropy-maximized configuration.) Note that we do not assume a priori that a black hole has a horizon or that the maximum entropy is given by the Bekenstein-Hawking formula.
@@
Solving the saturation condition for the entropy bound under the maximum excitation and using the consistency with local thermodynamics, we find uniquely the entropy-maximizing configuration for a fixed surface area ${\cal A}$ (Sec.IV). The interior metric $g_{\mu\nu}^{*}$ is given by
@@
We call it a semi-classical gravity condensate. Here, the near-Planckian curvature inside induces quantum fluctuations of various modes, which generate a large tangential pressure $\langle T^{\theta}{}_{\theta}\rangle_{*}$ associated with the 4D Weyl anomaly, making the condensate locally anisotropic. This pressure supports the system against the strong self-gravity, exceeds the Buchdahl limit Buchdahl , and self-consistently keeps the curvature finite. As a result, the energy is distributed throughout the interior, and the small central region ($0\leq r\lesssim\sqrt{\sigma}$) beyond the description by (1) has only a small energy and can be assumed to be almost flat. Thus, no singularity appears.
@@
This is close to but still outside $r=a_{0}$, and therefore, the configuration has no horizon but looks like a classical black hole from the outside. Indeed, it is almost black due to the exponentially large redshift of (3) CY .
We can also obtain the gravity condensate by considering a formation process. Generically, a configuration obtained by a thermodynamically reversible process should have the maximum entropy. In Refs.KMY ; KY1 ; KY4 ; KY5 , we considered a process in which thermal radiation comes together reversibly due to self-gravity in a heat bath at Hawking temperature, solved the self-consistent time evolution including the backreaction from Hawking-like radiation during the process, and obtained the metric (3). Therefore, the gravity condensate should be the most typical configuration with the maximum entropy according to the second law of thermodynamics, which is consistent with the above construction based on local typicality.
Now, to evaluate explicitly $S_{max}\equiv S[g_{\mu\nu}^{*};R)$, we apply the Unruh effect (or the local temperature due to the particle creation inside) and thermodynamic relations locally to the interior metric (3) and obtain the entropy density $s_{*}(r)$ (Sec.V). We evaluate (2) and derive
@@
where ${\cal A}\equiv 4\pi R^{2}=4\pi a_{0}^{2}+\mathcal{O}(1)$ for (4). Therefore, the maximum entropy $S_{max}$ coincides with the Bekenstein-Hawking formula, including the coefficient $1/4$. Here, the self-gravity changes the entropy (2) from the volume law to the area law Y1 . We then check the Hawking temperature and the negative specific heat.
This derives the Bousso bound for thermodynamic entropy (Sec.VI). Therefore, (5) means that the gravity condensate is derived as the unique configuration saturating the Bousso bound in our class of configurations. Furthermore, the interior metric (3) saturates the local sufficient conditions for the Bousso bound proposed in Refs.FMW ; BFM and should provide a clue to bulk dynamics rooted in the essence of holography.
We finally discuss 10 prospects for this picture of black holes (Sec.VII): role of self-gravity in holography, relation to other gravity-condensate models Dvali3 ; Oriti1 , description as a thermodynamic phase, path-integral evaluation of $S[g_{\mu\nu};R)$ BY1 ; BY2 ; Jacobson_vol , thermodynamic entropy vs entanglement entropy Minic ; Casini ; Jacobson_entangle , recovery of state-dependence in Hawking radiation KY2 , non-typical configurations, relation to the classical picture of black holes, gravitational field with finite entropy Jacobson ; Pad , and phenomenology echo ; CY . We then reach a speculative view that the gravity condensate represents a quantum-gravitational condensation in a semi-classical manner.
@@ II.2 Local typicality
In general, the order of the magnitude of the entropy in a system with a small volume $\Delta V$ and a high temperature $T$ can be estimated as $\sim\langle-T^{t}{}_{t}\rangle_{th}\Delta V/T$. For example, in a spherical ultra-relativistic fluid with volume $4\pi r^{2}\Delta\hat{r}$ (including self-gravity), the Stefan-Boltzmann law holds approximately Landau_SM ; fluidbook , and the entropy is given by $s(r)\Delta\hat{r}=\frac{4\langle-T^{t}{}_{t}(r)\rangle_{th}}{3T_{loc}(r)}4\pi r^{2}\Delta\hat{r}$ (see (77)). Note here that the non-locality of entropy BFM ; BCFM is considered for the width (8). Therefore, using this, (9) and (10), the order of the magnitude of the entropy in our subsystem can be estimated by $s(r)\Delta\hat{r}\sim\frac{\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle}{\epsilon(r)}4\pi r^{2}\Delta\hat{r}$, leading to [^8] One might think that in a pure state, entropy is zero. Microscopically, thermodynamic entropy is given by the number of possible states consistent with fixed macroscopic parameters (most such states are typical), which is state-independent Landau_SM . Physically, for example, a cold atomic system in a pure state will develop thermal behavior after a quench process, such that macroscopic quantities such as energy density have the same value as their thermal expectation values. Thus, assuming local typicality, the typical (i.e., thermal) behavior of the energy density can be used to estimate the entropy density in this phenomenological method. (See Sec.VII for microscopic methods.)
@@ II.3 Self-gravity
which provides the entropy $S[g_{\mu\nu};R)$ for the self-consistent solution to (1). Thus, the Hamiltonian constraint (13) relates one metric $g_{\mu\nu}$ and a set of the typical states $\{|\psi\rangle\}$ that have the same thermodynamic energy density $\langle-T^{t}{}_{t}(r)\rangle_{th}$, and the finite entropy (16) is obtained. [^9] Here are some comments on the subtleties of the above entropy evaluation.We use only the geometrical static condition and local consistencies with thermodynamics, and do not assume a priori a global thermodynamic equilibrium. This is motivated by two facts. More physically, even for static configurations, one should consider their formation processes. Generically, mechanical and global/local thermodynamic equilibria are different due to differences in relaxation timescales Landau_SM . In particular, we are now considering self-gravity. Therefore, it is necessary to consider time-delay effects in the formation process and discuss in which equilibrium state the configuration is for the timescale under consideration. The other is that, a notion of global thermodynamic equilibrium in self-gravitating systems, leading often to thermodynamical instability from negative specific heat Antonov ; Lynden , is still controversial Landau_SM ; Pad_thermo , and in particular, one consistent with (1) is not known. Furthermore, we do not assume a condition of isotropic fluid Green ; Xia (like thermal radiation Sorkin ) because it is not clear a priori whether such fluid maximizes entropy including self-gravity for a fixed surface area. Thus, we have estimated $S[g_{\mu\nu};R)$, (16), by using only the geometric static condition and the local consistency with thermodynamics. For the case of $S_{max}$, we will check the validity of this treatment by constructing the self-consistent solution $(g_{\mu\nu}^{*},|\psi\rangle_{*})$ (Sec.IV.2). In Sec.VII we will discuss a global equilibrium in quantum gravity briefly. In Appendix D we will argue that Tolman’s law does not lead to maximum entropy.
@@ III Upper bound
in the interior metric of (7) Sorkin . See Fig.4.
Figure 4: Meaning of the semi-classical static condition (18).
This means that, considering the non-locality of quanta, a quantum that constitutes a part at $r$ of a static configuration must be located at least its wavelength away from the Schwarzschild radius $a(r)$ of the energy inside it; otherwise, the trapped region would appear, the quantum would lie inside it and fall toward the center, and the configuration would not be static. (The validity of (18) will be discussed in Sec.IV.3.1.)
Using (11) and $\lambda\sim\frac{\hbar}{\epsilon}$, we can rewrite (18) as
@@
where we have used $4\pi\int^{R}_{0}drr^{3}\langle-T^{t}{}_{t}(r)\rangle\leq 4\pi R\int^{R}_{0}drr^{2}\langle-T^{t}{}_{t}(r)\rangle=R\frac{a(R)}{2G}\approx R\frac{a_{0}}{2G}\equiv RM_{0}$ from (13). [^10] In Ref.Sorkin , (III) and (22) were obtained only for thermal radiation, while our derivation can be applied to more general cases. Note that $G$ does not appear here, but this is a result of the dynamics of gravity because we need $\mathcal{H}=0$ (13) to obtain the ADM energy $M_{0}$.
@@ IV.1 Saturating energy distribution $a_{*}(r)$
The estimation (16) and bound (17) mean that the maximum entropy can be obtained by setting $\epsilon_{*}(r)=\epsilon_{max}$ at each $r$. Therefore, (23) becomes $nl_{p}^{2}\sim r(r-a_{*}(r))$, leading to
@@
with $\sigma=fnl_{p}^{2}$ ($f$: a dimensionless positive constant of $\mathcal{O}(1)$). Thus, the maximum entropy $S_{max}\equiv S[g_{\mu\nu}^{*},R)$ is estimated from (21) and (24) as
@@
where $a_{0}\approx a_{*}(R)\approx R$. (In Sec.V.2, the precise relation of $a_{0}$ and $R$ will be fixed.) Note that the saturation of both (17) and (18) corresponds to the strong-gravity limit mentioned in Sec.I.
@@
where we have applied $\epsilon(r)=\epsilon_{max}$ and (24) to (14). This means that $\mathcal{O}(\sqrt{n})$ bit of information is packed per the Planck length constantly KY2 ; Y1 . Thus, a necessary condition for the entropy maximization is the radial uniformity. Conversely, as a result of the radial uniformity, (24) can be uniquely obtained as a solution satisfying the entropy maximization condition (23), without using $\epsilon(r)=\epsilon_{max}$ (see Appendix C). Therefore, from these two arguments, a necessary and sufficient condition for the entropy maximization is the radial uniformity. This is a non-trivial result for a self-gravitating system.
@@
On the other hand, using (13) and (24), we have the energy density $\langle-T^{t}{}_{t}\rangle_{*}\approx\frac{1}{8\pi Gr^{2}}$ and the local energy in each subsystem, $\Delta E_{loc}^{*}(r)\equiv 4\pi r^{2}\langle-T^{t}{}_{t}\rangle_{*}\Delta\hat{r}_{*}\sim\sqrt{n}m_{p}$. Therefore, we must have at least $\frac{a_{*}(r)}{2G}\gtrsim\sqrt{n}m_{p}$, meaning that (24) holds only in